solve for EBIT ((EBIT - 919,000,000)*(1-.35))/1244000000 = ((EBIT - 552,000,000)*(1-.35))/929,000,000

solve for EBIT

((EBIT - 919,000,000)*(1-.35))/1244000000 = ((EBIT - 552,000,000)*(1-.35))/929,000,000
Answer #1

((EBIT - 919,000,000) * (1-.35))/1244000000 = ((EBIT - 552,000,000) * (1-.35))/929,000,000

Divide the equation by 1 million.

((EBIT - 919) *(1-.35))/1244 = ((EBIT - 552) * (1-.35))/929

((EBIT - 919) *(1-.35)) * 929 = ((EBIT - 552) * (1-.35)) * 1244

(EBIT - 919) * 603.85 = (EBIT – 552) * 808.6

603.85EBIT – (919 * 603.85) = 808.6EBIT – (552 * 808.6)

603.85EBIT – 554,938.15 = 808.6EBIT – 446,347.2

808.6EBIT – 603.85EBIT = -554,938.15 + 446,347.2

204.75EBIT = -108,590.95

EBIT = -108,590.95 / 204.75

EBIT = -530.35873

Multiply by 1 million

EBIT = -530.35873 * 1,000,000 = -530,358,730

EBIT = -530,358,730